OpenStudy (anonymous):

2. For her chemistry experiment, Marisol needs 2 liters of a 50% alcohol solution. The lab has two containers, one with 40% alcohol solution and the other with 80% alcohol solution. i cant figure out the linear equations

OpenStudy (anonymous):

x+y=2 given 40% alcohol and 80% alcohol solution we would then have .4x+.8y=1

OpenStudy (anonymous):

ok thank you i had that same answer i just wasn't sure if it was right :)

OpenStudy (anonymous):

Anytime!

OpenStudy (anonymous):

so would x and y both equal 1? @reeses

OpenStudy (anonymous):

its a system of equations so solve for x for one equation and plug it into x for the other equation then the value get for x you plug it back into the first equation then solve for y :)

OpenStudy (victoriasushchik):

Let the amount of 40% solution be a and 80% solution b. so, 0.4a + 0.8b = 0.5(a + b) Now, a + b = 2 => 0.4a + 0.8(2 - a) = 0.5(2) i.e. 1.6 - 0.4a = 1 so, 0.4a = 0.6 => a = 1.5 and b = 0.5 Hence, 1.5 litres of 40% solution needs mixing with 0.5 litres of 80% solution to give a 2-litre, 50% solution :)>

OpenStudy (denisaboichuk):

1.5 and 0.5